OpenStudy (anonymous):

Using the point slope form y-ybase1 =m(x-xbase1), write an equation for the line 1. with slope -2, containing the point (3,4) 2. containing the points (1,-3) and (-5,2) 3. with slope 0, containing the point (4,2) 4. perpendicular to the line in prob. #1, containing the point (3,4)

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

1) put in formula \[y-4=-2(x-3)\] distribute \[y-4=-2x+6\] add 4 \[y=-2x+10\]

OpenStudy (anonymous):

2) slope is \[m=\frac{-3-2}{1+5}=\frac{-5}{6}\] now repeat steps above